When a merger takes place between two companies to form a single firm, the target company to operate as a separate identity. Consider the following scenario: Universal Drapers Inc. is considering an acquisition of Mammoth Pictures Inc, and estimates that acquiring Mammoth will result in incremental after-tax net cash flows in years 1-3 of $17.0 million, $25.5 million, and $30.6 million, respectively. After the first three years, the incremental cash flows contributed by the Mammoth acquisition are expected to grow at a constant rate of 4% per year, Universals current beta is 0.40, but its post-merger beta is expected to be 0.52. The risk-free rate is 5.5%, and the market risk premium is 7.60%. Based on this information, complete the folowng table by selecting the appropriate values: Value Post-merger cost of equity Projected value of the cash flows at the end of three years The value of Mammoth Pictures Inc.s contribution to Universal Drapers Inc. Mammoth Pictures Inc. has 6 million shares of common stock outstanding. What is the largest tender offer Universal Drapers Inc. should makeon each of Mammoth Pictures Inc.s shares? O $84.24 O $67.40 O $101.09
Answer #1

1)

When a merger takes place between two companies to form a single firm, the target company does not continue to operate as a separate identity

it can also be said as

When a merger takes place between two companies to form a single firm, the target company ceases to operate as a separate identity

2)

post merger cost of equity (r)=risk free rate+post-merger beta*market risk premium

=5.5%+0.52*7.60%

=9.4520% or 9.45%

projected value of the cash flows at the end of three years=(CF3*(1+g))/(r-g)

=(30.6*(1+4%))/(9.452%-4%)

= 583.7124 million or 583.71 million

the value of Mammoth Pictures's contribution to Universal=CF1/(1+r)^1+CF2/(1+r)^2+CF3/(1+r)^3+((CF3*(1+g))/(r-g))/(1+r)^3

=17/(1+9.452%)^1+25.5/(1+9.452%)^2+30.6/(1+9.452%)^3+((30.6*(1+4%))/(9.452%-4%))/(1+9.452%)^3

= 505.33 million

largest tender offer per share=505.33/6=84.2 or 84.24

the above is answer..